Slippery Slope Roof Contracting has an equity beta of 1.2, capital structure with 2/3 debt, and a zero tax rate. What is their asset beta? 
Answer #1

Debt=2/3

Hence equity will be =(1-2/3)=1/3

Hence debt/equity=(2/3)/(1/3)=2

Asset beta=Beta of equity/[1+(1-tax)*debt /equity ]

= 1.2/[1+(1-0)*2/]

=1.2/3

=0.4.